OpenStudy (anonymous):

If l = 12 cm and m = 5 cm, what is the length of n?

OpenStudy (calculusxy):

Is there an equation that is attached to this?

OpenStudy (anonymous):

its a triangle

OpenStudy (anonymous):

If l = 12 cm and m = 5 cm, what is the length of n?

OpenStudy (calculusxy):

You have to use the Pythagorean Theorem. Follow what I do, and if you have any questions be sure to tell me. a^2+b^2=c^2 This is the formula. But don't get worried by the variables, you can always switch a and b for l and m. Also c can be switched for n. So not to get this confusing, I will substitute the letterings for you, l^2+m^2=n^2 Lets now substitute the variables with the numbers. 12^2+5^2=c^2 169=c^2 Remember that n equals to the square root of 169. n= √169 n=13

OpenStudy (calculusxy):

@kthy1347 Do you get it?